OpenStudy (anonymous):

Write the equation of the circle that has a diameter with endpoints (6,-1) and (0,-7).

OpenStudy (jdoe0001):

hmmm  what's the radius of it anyway?

OpenStudy (anonymous):

This is the whole question, we have to figure out the radius to make the question.

OpenStudy (jdoe0001):

anyway, the raidus will be half the diameter to get the diameter, get the distance between the endpoints that is \(\bf \textit{distance between 2 points}\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ % (a,b) &({\color{red}{ 6}}\quad ,&{\color{blue}{ -1}})\quad % (c,d) &({\color{red}{ 0}}\quad ,&{\color{blue}{ -7}}) \end{array}\qquad % distance value d = \sqrt{({\color{red}{ x_2}}-{\color{red}{ x_1}})^2 + ({\color{blue}{ y_2}}-{\color{blue}{ y_1}})^2}\) and then, find the center of the circle which happens to be at at the midpoint of the those 2 endpoints that is \(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ % (a,b) &({\color{red}{ 6}}\quad ,&{\color{blue}{ -1}})\quad % (c,d) &({\color{red}{ 0}}\quad ,&{\color{blue}{ -7}}) \end{array}\qquad % distance value \left(\cfrac{{\color{red}{ x_2}} + {\color{red}{ x_1}}}{2}\quad ,\quad \cfrac{{\color{blue}{ y_2}} + {\color{blue}{ y_1}}}{2} \right)\)

OpenStudy (jdoe0001):

and once you find the diameter half of that is the radius and then use that with the center for the circle's equation \(\bf (x-{\color{brown}{ h}})^2+(y-{\color{blue}{ k}})^2={\color{purple}{ r}}^2 \qquad center\ ({\color{brown}{ h}},{\color{blue}{ k}})\qquad radius={\color{purple}{ r}}\)

OpenStudy (anonymous):

Thanks!

OpenStudy (jdoe0001):

yw

OpenStudy (jdoe0001):

I got a radical for the diameter btw